# Fast Adaptive Anti-Jamming Channel Access via Deep Q Learning and Coarse-Grained Spectrum Prediction

### v2

@@ Abstract
This paper investigates the anti-jamming channel access problem in complex and unknown jamming environments, where the jammer could dynamically adjust its strategies to target different channels. Traditional channel hopping anti-jamming approaches using fixed patterns are ineffective against such dynamic jamming attacks. Although the emerging deep reinforcement learning (DRL) based dynamic channel access approach could achieve the Nash equilibrium under fast-changing jamming attacks, it requires extensive training episodes. To address this issue, we propose a fast adaptive anti-jamming channel access approach guided by the intuition of “learning faster than the jammer", where a synchronously updated coarse-grained spectrum prediction serves as an auxiliary task for the deep Q learning (DQN) based anti-jamming model. This helps the model identify a superior Q-function compared to standard DRL while significantly reducing the number of training episodes. Numerical results indicate that the proposed approach significantly accelerates the rate of convergence in model training, reducing the required training episodes by up to 70% compared to standard DRL. Additionally, it also achieves a 10% improvement in throughput over NE strategies, owing to the effective use of coarse-grained spectrum prediction.
Keywords Anti-jamming  $\cdot$ fast adaptive channel access  $\cdot$ coarse-grained spectrum prediction  $\cdot$ deep Q learning reinforcement learning
@@ 1 Introduction
Wireless communications have found extensive applications in both civilian and military scenarios, where ensuring robust anti-jamming capabilities is paramount for secure transmissions, given the vulnerability of wireless links Wang et al. (2020); Tang et al. (2019). The dynamic characteristics of wireless channels present significant challenges for conventional model-based anti-jamming methods Xiao et al. (2018); Lv et al. (2023); Jia et al. (2019). It has been shown that the model-free approaches, such as reinforcement learning (RL) Watkins and Dayan (1992); Mnih et al. (2015), could help to empower the legitimate user with enhanced anti-jamming capability in wireless communication environments that are dynamic and unknown Liu et al. (2018); Xiao et al. (2020). However, jammers have also evolved intelligently with advancements in Universal Software Radio Peripheral (USRP) Zhu et al. (2016) and artificial intelligence (AI) Amuru et al. (2016) in recent years. Significant breakthroughs have recently been made in the field of smart jamming and intelligent jamming Wang et al. (2019); Feng et al. (2022); Qi et al. (2021); Bout et al. (2023); Wang et al. (2022).
@@
Considerable efforts have been devoted to combating intelligent jamming attacks. In Xiao et al. (2021), the authors treated the jammer as an integral component of the environment and introduced a DRL-based algorithm for combating an RL-based jammer. Considering the non-stationarity characteristic of the environment where the jammer could also be regarded as an RL agent Pirayesh and Zeng (2022), several works have formulated the interaction between the legitimate user and the intelligent jammer within a game theory framework. Xiao Xiao et al. (2018) proposed a Two-Dimensional anti-jamming communication scheme using a hotbooting deep Q-network to enhance mobile device utility and signal quality under cooperative jamming attacks. In Zhang and Wu (2024), Zhang explored the interactions between the legitimate user with faking-slot transmission and the intelligent reactive jammer within a bi-matrix game framework and derived the equilibrium for the game using the quadratic programming method. Additionally, a Neural Fictitious Self-Play (NFSP) method was proposed in Li et al. (2022) to identify the approximate Nash equilibrium (NE) solution for the dynamic game of radar anti-jamming with imperfect information by employing Minimax Q network Zhu and Zhao (2022), and a deceptively adversarial attack approach was proposed in Li et al. (2021) to tackle the challenge presented by smart jamming.
@@
On the other hand, modeling the interaction between players using multi-agent reinforcement learning and leveraging the behaviors of other agents in multi-agent systems (MAS) has gained considerable attention in recent years He et al. (2016); Lowe et al. (2017); Hong et al. (2018); Lanctot et al. (2017); Lu et al. (2022); Yu et al. (2022). The use of deep reinforcement opponent network (DRON) He et al. (2016) could achieve superior performance over that of DQN and its variants in multi-agent environments. Moreover, the model-based opponent modeling (MBOM) Yu et al. (2022) could simulate the iterative reasoning process within the environment model and generate a range of opponent policies for achieving a more effective adaptation in a variety of tasks. To address the anti-jamming problems in wireless communications, Li Li et al. (2023a) proposed an opponent modeling based anti-intelligent jamming (OMAIJ) algorithm that analyzes the jammer’s policy and targets its vulnerabilities. Yuan Yuan et al. (2024) introduced an opponent awareness-based anti-jamming algorithm that considers the jammer’s learning to effectively counter intelligent jamming attacks.
@@
In this paper, a novel fast adaptive channel access approach, which combines DQN and coarse-grained spectrum prediction, is proposed for anti-jamming. The proposed approach demonstrates a faster convergence rate than both the DRL-based approach and the opponent modeling approach in model training, outperforming the Nash equilibrium in scenarios involving DRL-based jammers. The contributions of this paper are outlined below.
  • Firstly, we describe the adversarial scenario in which a fixed-mode jammer and a DRL-based jammer operate simultaneously. The interaction between the legitimate user and the DRL-based jammer could be formulated as a Markov Game, where the user and the jammer have completely opposing objectives.
@@
The remainder of this paper is organized as follows. The interaction between the legitimate user and the jammers is formulated in Section II. Subsequently, a supervised learning based coarse-grained spectrum prediction scheme is introduced in Section III, followed by a novel fast adaptive anti-jamming channel access approach with joint DQN and coarse-grained spectrum prediction in Section IV. Simulation results regarding the anti-jamming performance and convergence rate in model training of the proposed approach are provided in Section V. Section VI concludes the paper.
@@ 2.1 System Model
Assuming that the communication band $\left[f_{L},f_{U}\right]$, with total bandwidth $B=f_{U}-f_{L}$, could be divided into $M\in\mathbb{Z}^{+}$ non-overlapping channels. The available channel set is denoted as =$\mathcal{F}=\left\{f_{1},f_{2},\ldots,f_{M}\right\}$, where the bandwidth of each channel is $b=B/M$. We consider a synchronous time-slotted system, in which the basic time slots of the legitimate user and the DRL-based intelligent jammer are perfectly aligned. The duration of each basic time slot is $\Delta t$, which is the smallest unit of time resolution. For convenience, the term “basic time slot” is abbreviated as “time slot" in what follows. During the $k$-th time slot (i.e., from $(k-1)\Delta t$ to $k\Delta t$), the legitimate user selects an available channel $f_{k}^{u}\in\mathcal{F}$ for transmission, while the intelligent jammer targets $N_{I}$ consecutive channels in $\mathcal{I}_{k}=\left\{f_{k}^{i}\right\}_{i=1,2,\ldots,N_{I}}$ and the fixed-mode jammer releases jamming signals on channel $f_{k}^{s}$. Similar to Wu et al. (2013), the block fading channel model is assumed in the proposed anti-jamming model. Specifically, the channel gain from the transmitter to the receiver during the $k$-th time slot is defined as
@@
The illustration of the spectrum vectors and the coarse-grained spectrums in different hops are shown in Fig. 2 and Fig. 3, respectively. The small rectangles in Fig. 2 represent the samples of spectrum vectors, and the rectangles in Fig. 3 represent the samples of coarse-grained spectrums. The time resolution and the frequency resolution of spectrum analysis are $\Delta t=T_{h}/N_{h}$ and $\Delta f=B/N_{F}$, respectively, and the spectrum defined in (5) could be employed to learn the behavior of both the legitimate user and the jammers during each time slot. While the time resolution and the frequency resolution of coarse-grained spectrum analysis are $T_{h}$ and $\Delta f^{\prime}=B/M=b$, respectively, and the coarse-grained spectrum defined in (7) is used to express the channel state during each hop with $M$ samples. It is noted that, when representing the spectrum within the same time duration and frequency range, the number of samples in the spectrum matrix $\left[{\bf{s}}_{(n-1)N_{h}+1},\ldots,{\bf{s}}_{nN_{h}-1},{\bf{s}}_{nN_{h}}\right]^{\top}$ is $(N_{h}\times\frac{N_{F}}{M})$ times greater than the coarse-grained spectrum ${\bf{c}}_{n}$.
The Signal-to-Interference-plus-Noise Ratio (SINR) at the legitimate receiver defined in (9) is often utilized to evaluate the quality of the received signal. If the SINR exceeds the given demodulation threshold $\beta_{\text{th}}$, the user successfully mitigates jamming attacks. Otherwise, the legitimate transmission fails. The legitimate user aims to find an appropriate channel at the beginning of each hop for achieving a receiving SINR that exceeds $\beta_{\text{th}}$ during each time slot of the hop.
[…]
2.2 Markov Game Model
In the proposed anti-jamming scenario, the adaptive channel access decision-making process of the legitimate user is sequential, and the nonstationarity evolution of the state happens when the intelligent jammer adopts a time-varying jamming policy. Under these circumstances, the evolution of the environment is influenced by the actions of both the legitimate transmitter and the intelligent jammer, rather than solely by the action of the legitimate transmitter. This non-stationary characteristic prompts us to model the interaction between the legitimate user and the intelligent jammer as a Markov Game (MG), where the user and the jammer could make decisions simultaneously, each pursuing completely opposing objectives. The anti-jamming MG could be described by a seven-tuple, namely,
where $\mathcal{S}$ represents the set of environment states, $\mathcal{A}^{u}$ and $\mathcal{A}^{j}$ represent the action set of the legitimate user and the intelligent jammer, respectively, $\mathcal{P}$ is the transition function, $\mathcal{R}^{u}$ denotes the legitimate user’s reward function, $\mathcal{R}^{j}$ denotes the intelligent jammer’s reward function, and $\gamma$ is the discount factor.
For the considered anti-jamming MG, we assume that the actions of both the legitimate user and the intelligent jammer are allowed to change among hops. An illustrative diagram of the time slot structure for the proposed anti-jamming MG is illustrated in Fig. 2. Since the spectrum waterfall defined in Chen and Wen (2016) contains time, frequency and power domain information, it could be used to represent the complex spectrum state and provide enough information for either anti-jamming or jamming decision-making. Therefore, the environment state of the $n$-th hop could be represented as the observed spectrum waterfall at the beginning of the $n$-th hop, i.e.,
where ${\bf{s}}_{i}$ denotes the spectrum vector during the $i$-th time slot, $N_{T}$ denotes the length of historical data. The user takes ${\bf{S}}_{n}$ as the input for anti-jamming decision-making. Similarly, the DRL-based intelligent jammer also takes the observed spectrum waterfall ${\bf{S}}_{n}^{j}$ as the input for its jamming decisions. For convenience, ${\bf{S}}_{n}^{j}={\bf{S}}_{n}$ is used in what follows.
Let $\pi(\cdot)$ and $\mu(\cdot)$ denote the policies of the legitimate user and intelligent jammer, respectively. During the $n$-th hop, the legitimate user executes an anti-jamming action $a_{n}^{u}\in\mathcal{F}$ based on the policy $\pi({\bf{S}}_{n})$, while the intelligent jammer performs a jamming action $a_{n}^{j}\subseteq\mathcal{F}$ according to the policy $\mu({\bf{S}}_{n})$. Then, the state ${\bf{S}}_{n}$ transits to the next state ${\bf{S}}_{n+1}$ with probability $\mathcal{P}({\bf{S}}_{n+1}|{\bf{S}}_{n},a_{n}^{u},a_{n}^{j})$, and the environment provides immediate rewards $r_{n}^{u}=\mathcal{R}^{u}({\bf{S}}_{n},a_{n}^{u},a_{n}^{j})$ and $r_{n}^{j}=\mathcal{R}^{j}({\bf{S}}_{n},a_{n}^{u},a_{n}^{j})$ to the legitimate user and the intelligent jammer, respectively.
At the end of each hop, the receiver responds to the transmitter with a feedback (i.e., ACK/NACK) through the control link. Specifically, if $\beta(k,f_{u}^{k})\geq\beta_{\text{th}},\ \forall k=(n-1)N_{h}+1,\ldots,nN_{h}-1,nN_{h}$ holds during the $n$-th hop, the receiver transmits an ACK signal to the transmitter. Otherwise, an NACK signal is transmitted. Motivated by Yuan et al. (2024), we assume that the legitimate user’s agent is located at the receiver. The agent could determine which channel to access and transmit the channel decision message $a_{n}^{u}$ to the transmitter at the beginning of the $n$-th hop through the control link. Then it takes the minimum SINR during the $n$-th hop as the immediate reward for action $a_{n}^{u}$, namely,
Whereas the intelligent jammer has a completely contrasting objective. It is assumed that the worst jammer could accurately assess the effectiveness of jamming attacks, and the reward for the intelligent jammer’s action $a_{j}^{n}$ is determined by the intercepted ACK/NACK signal from the control link during the $n$-th hop, namely,
$r^{j}_{n}=\begin{cases}1,\quad&\text{if the NACK feedback is detected},\\ -1,\quad&\text{if the ACK feedback is detected}.\end{cases}$ (13)
@@ 2.2 Markov Game Model
In the process of the anti-jamming game, the legitimate user aims to maximize its cumulative reward $R^{u}({\bf{S}}_{1})$, i.e.,
@@
Meanwhile, the intelligent jammer aims to maximize its cumulative reward $R^{j}({\bf{S}}_{1})$, i.e.,
@@
(BR in the anti-jamming MG) Given the policy of the intelligent jammer $\mu$, the policy $\pi^{b}$ of the legitimate user is defined as the BR policy if there exists no alternative policy that could yield a higher cumulative reward, formally expressed as
@@
Conversely, when the legitimate user employs the policy $\pi$, the BR policy of the intelligent jammer, denoted as $\mu^{b}$, must satisfy
@@
(NE in the anti-jamming MG) The NE is defined as a pair of policies $(\pi^{*},\mu^{*})$, where both $\pi^{*}$ and $\mu^{*}$ represent the best responses to each other. Formally, this could be expressed as
@@
According to Nash (1951), the NE of the proposed anti-jamming MG always exists and is equivalent to the minimax solution of the game, i.e.,
@@
The NE delineates the maximum return that a legitimate user could achieve when facing a formidable opponent. It is particularly meaningful when the jammer is capable of adapting its policy in response to the user’s actions. When the legitimate user adopts the policy $\pi^{*}$, its return is guaranteed to be at least equal to the Nash equilibrium. While, if the legitimate user switches to an alternative anti-jamming policy, its return might fall below $R^{u}({\bf{S}}_{1};\pi^{*},\mu^{*})$.
Minimax Q learning Wang et al. (2011) could be employed to determine the NE for MGs without any prior knowledge of the environment dynamics. Additionally, the legitimate user could achieve a beyond NE performance by employing an opponent modeling based DQN Li et al. (2023a). However, these methods assume that the legitimate user could access the actions selected by the intelligent jammer at the beginning of each hop, which is generally impractical. Meanwhile, these DRL-based methods necessitate extensive training episodes in non-stationary environments, where the strategies of both the legitimate user and the intelligent jammer could continuously evolve over time. To address these issues, we propose a fast adaptive anti-jamming channel access approach to find a policy beyond NE with fewer training episodes in the following sections.

### v3

@@ Abstract
This paper investigates the anti-jamming channel access problem in complex and unknown jamming environments, where the jammer could dynamically adjust its strategies to target different channels. Traditional channel hopping anti-jamming approaches using fixed patterns are ineffective against such dynamic jamming attacks. Although the emerging deep reinforcement learning (DRL) based dynamic channel access approach could achieve the Nash equilibrium (NE) under fast-changing jamming attacks, it requires extensive training episodes. To address this issue, we propose a fast adaptive anti-jamming channel access approach guided by the intuition of “learning faster than the jammer", where a synchronously updated coarse-grained spectrum prediction serves as an auxiliary task for the deep Q network (DQN) based anti-jamming model. This helps the model identify a superior Q-function compared to standard DRL while significantly reducing the number of training episodes. Numerical results indicate that the proposed approach significantly accelerates the rate of convergence in model training, reducing the required training episodes by up to 70% compared to standard DRL. Additionally, it also achieves a 10% improvement in throughput over NE strategies, owing to the effective use of coarse-grained spectrum prediction.
Keywords Anti-jamming $\cdot$ fast adaptive channel access $\cdot$ coarse-grained spectrum prediction $\cdot$ deep Q learning
@@ 1 Introduction
Wireless communications have found extensive applications in both civilian and military scenarios, where ensuring robust anti-jamming capabilities is paramount for secure transmissions, given the vulnerability of wireless links Wang et al. (2020); Tang et al. (2019). The dynamic characteristics of wireless channels present significant challenges for conventional model-based anti-jamming methods Xiao et al. (2018a); Lv et al. (2023); Jia et al. (2019). It has been shown that the model-free approaches, such as reinforcement learning (RL) Watkins and Dayan (1992); Mnih et al. (2015), could help to empower the legitimate user with enhanced anti-jamming capability in wireless communication environments that are dynamic and unknown Liu et al. (2018); Xiao et al. (2020). However, jammers have also evolved intelligently with advancements in Universal Software Radio Peripheral (USRP) Zhu et al. (2016) and artificial intelligence (AI) Amuru et al. (2016) in recent years. Significant breakthroughs have recently been made in the field of smart jamming and intelligent jamming Wang et al. (2019); Feng et al. (2022); Qi et al. (2021); Bout et al. (2023); Wang et al. (2022).
@@
Considerable efforts have been devoted to combating intelligent jamming attacks. In Xiao et al. (2021), the authors treated the jammer as an integral component of the environment and introduced a DRL based algorithm for combating an RL-based jammer. Xu Xu et al. (2020) proposed a Deep Recurrent Q-Network (DRQN) based channel access method to identify an optimal access policy that minimizes conflict probability and simultaneously maximizes channel utilization in multi-user scenarios. Additionally, the DRL-based algorithm has potential applications in underwater anti-jamming communications Xiao et al. (2018b) and unmanned aerial vehicles (UAV) anti-jamming communications Wang et al. (2023). Li Li et al. (2023a) proposed a Markov decision process (MDP) with a two-dimensional action space (transmit frequency and power) and introduced a dual action network based DRL algorithm with an action feedback mechanism to address the anti-jamming problem in frequency hopping (FH) communication systems. A parallel policy network-based DRL algorithm, which adapts transmission power while accessing idle channels, was introduced in Li et al. (2022a) to address the anti-jamming MDP problem. In Li et al. (2023b), a labeled DRL-based dynamic spectrum anti-jamming approach was proposed to adapt to the fast-changing jamming environment. Considering the non-stationarity characteristic of the environment where the jammer could also be regarded as an RL agent Pirayesh and Zeng (2022), several works have formulated the interaction between the legitimate user and the intelligent jammer within a game theory framework. Xiao Xiao et al. (2018a) proposed a Two-Dimensional anti-jamming communication scheme using a hotbooting deep Q-network to enhance mobile device utility and signal quality under cooperative jamming attacks. In Zhang and Wu (2024), Zhang explored the interactions between the legitimate user with faking-slot transmission and the intelligent reactive jammer within a bi-matrix game framework and derived the equilibrium for the game using the quadratic programming method. Additionally, a Neural Fictitious Self-Play (NFSP) method was proposed in Li et al. (2022b) to identify the approximate NE solution for the dynamic game of radar anti-jamming with imperfect information by employing Minimax Q network Zhu and Zhao (2022), and a deceptively adversarial attack approach was proposed in Li et al. (2021) to tackle the challenge presented by smart jamming.
@@
On the other hand, modeling the interaction between players using multi-agent reinforcement learning and leveraging the behaviors of other agents in multi-agent systems has gained considerable attention in recent years He et al. (2016); Lowe et al. (2017); Hong et al. (2018); Lanctot et al. (2017); Lu et al. (2022); Yu et al. (2022). The use of deep reinforcement opponent network (DRON) He et al. (2016) could achieve superior performance over that of DQN and its variants in multi-agent environments. Moreover, the model-based opponent modeling (MBOM) Yu et al. (2022) could simulate the iterative reasoning process within the environment model and generate a range of opponent policies for achieving a more effective adaptation in a variety of tasks. To address the anti-jamming problems in wireless communications, Li Li et al. (2023c) proposed an opponent modeling based anti-intelligent jamming (OMAIJ) algorithm that analyzes the jammer’s policy and targets its vulnerabilities. Yuan Yuan et al. (2024) introduced an opponent awareness-based anti-jamming algorithm that considers the jammer’s learning to effectively counter intelligent jamming attacks.
@@
In this paper, a novel fast adaptive channel access approach, which combines DQN and coarse-grained spectrum prediction, is proposed for anti-jamming. Unlike previous works that focus exclusively on reinforcement learning or opponent modeling, we propose using coarse-grained spectrum prediction as an auxiliary task to accelerate the rate of convergence during the training of the DQN-based anti-jamming model and identify a superior Q-function compared to standard DRL. This integration of spectrum prediction with DQN represents a key innovation, enabling the model to identify fast adaptive anti-jamming strategies more effectively and efficiently, which is not addressed in prior studies. The proposed approach exhibits a faster rate of convergence than both the DRL-based approach and the opponent modeling approach during model training, outperforming the Nash equilibrium in scenarios involving DRL-based jammers. We believe these contributions provide a significant improvement over existing anti-jamming methods, addressing the limitations of the slow rate of convergence during training and the lack of adaptability to dynamic jamming strategies seen in earlier work. The contributions of this paper are outlined below.
  • Firstly, we describe the adversarial scenario in which a fixed-mode jammer and a DRL-based jammer operate simultaneously. The interaction between the legitimate user and the DRL-based jammer could be formulated as a Markov Game (MG), where the user and the jammer have completely opposing objectives.
@@
The remainder of this paper is organized as follows. The interaction between the legitimate user and the jammers is formulated in Section II. Subsequently, a supervised learning based coarse-grained spectrum prediction scheme is introduced in Section III, followed by a novel fast adaptive anti-jamming channel access approach with joint DQN and coarse-grained spectrum prediction in Section IV. Simulation results regarding the anti-jamming performance and the rate of convergence rate in model training of the proposed approach are provided in Section V. Section VI concludes the paper.
@@ 2.1 System Model
Assuming that the communication band $\left[f_{L},f_{U}\right]$, with total bandwidth $B=f_{U}-f_{L}$, could be divided into $M\in\mathbb{Z}^{+}$ non-overlapping channels. The available channel set is denoted as $\mathcal{F}=\left\{f_{1},f_{2},\ldots,f_{M}\right\}$, where the bandwidth of each channel is $b=B/M$. We consider a synchronous time-slotted system, in which the basic time slots of the legitimate user and the DRL-based intelligent jammer are perfectly aligned. The duration of each basic time slot is $\Delta t$, which is the smallest unit of time resolution. For convenience, the term “basic time slot” is abbreviated as “time slot" in what follows. During the $k$-th time slot (i.e., from $(k-1)\Delta t$ to $k\Delta t$), the legitimate user selects an available channel $f_{k}^{u}\in\mathcal{F}$ for transmission. Concurrently, the fixed-mode jammer releases jamming signals on channel $f_{k}^{s}$, while the intelligent jammer targets $N_{I}$ consecutive channels in $\mathcal{I}_{k}=\left\{f_{k}^{i}\right\}_{i=1,2,\ldots,N_{I}}$. This setting reflects the typical contiguous-band emission of practical wideband jammers and establishes a strong adversarial scenario for evaluating the robustness of anti-jamming strategies Pirayesh and Zeng (2022). Similar to Wu et al. (2013), the block fading channel model is assumed in the proposed anti-jamming model. Specifically, the channel gain from the transmitter to the receiver during the $k$-th time slot is defined as
@@
$c_{n}^{m}=\int_{(m-1)\Delta f^{\prime}}^{m\Delta f^{\prime}}S_{n}(f+f_{L})df,\quad m=1,2,\ldots,M,$ (8)
The illustration of the spectrum vectors and the coarse-grained spectra in different hops are shown in Fig. 2 and Fig. 3, respectively. The small rectangles in Fig. 2 represent the samples of spectrum vectors, and the rectangles in Fig. 3 represent the samples of coarse-grained spectra. The time resolution and the frequency resolution of spectrum analysis are $\Delta t=T_{h}/N_{h}$ and $\Delta f=B/N_{F}$, respectively. The spectrum defined in eq. (5) could be employed to learn the behavior of both the legitimate user and the jammers during each time slot. While the time resolution and the frequency resolution of coarse-grained spectrum analysis are $T_{h}$ and $\Delta f^{\prime}=B/M=b$, respectively. The coarse-grained spectrum defined in eq. (7) is used to express the channel state during each hop with $M$ samples. It is noted that, when representing the spectrum within the same time duration and frequency range, the number of samples in the spectrum matrix $\left[{\bf{s}}_{(n-1)N_{h}+1},\ldots,{\bf{s}}_{nN_{h}-1},{\bf{s}}_{nN_{h}}\right]^{\top}$ is $(N_{h}\times\frac{N_{F}}{M})$ times greater than the coarse-grained spectrum ${\bf{c}}_{n}$.
The Signal-to-Interference-plus-Noise Ratio (SINR) at the legitimate receiver defined in eq. (9) is often utilized to evaluate the quality of the received signal. If the SINR exceeds the given demodulation threshold $\beta_{\text{th}}$, the user successfully mitigates jamming attacks. Otherwise, the legitimate transmission fails. The legitimate user aims to find an appropriate channel at the beginning of each hop for achieving a receiving SINR that exceeds $\beta_{\text{th}}$ during each time slot of the hop.
$\beta(k,f_{u}^{k})=\frac{h_{k}^{u,r}(f_{k}^{u})\int_{-b/2}^{b/2}U(f)df}{\int_{f_{k}^{u}-b/2}^{f_{k}^{u}+b/2}\left[h_{k}^{s,r}(f_{k}^{s})J_{s}(f-f_{k}^{s})+\sum\limits_{i=1}^{N_{I}}h_{k}^{i,r}(f_{k}^{i})J_{i}(f-f_{k}^{i})+N_{k}(f)\right]df}.$ (9)
2.2 Markov Game Model
In the proposed anti-jamming scenario, the adaptive channel access decision-making process of the legitimate user is sequential, and the nonstationarity evolution of the state happens when the intelligent jammer adopts a time-varying jamming policy. Under these circumstances, the evolution of the environment is influenced by the actions of both the legitimate transmitter and the intelligent jammer, rather than solely by the action of the legitimate transmitter. This non-stationary characteristic prompts us to model the interaction between the legitimate user and the intelligent jammer as an MG, where the user and the jammer could make decisions simultaneously, each pursuing completely opposing objectives. The anti-jamming MG could be described by a tuple, namely,
where $\mathcal{S}$ represents the set of environment states, $\mathcal{A}^{u}$ and $\mathcal{A}^{j}$ represent the action set of the legitimate user and the intelligent jammer, respectively, $\mathcal{P}$ denotes the transition function, $\mathcal{R}^{u}$ is the legitimate user’s reward function, $\mathcal{R}^{j}$ is the intelligent jammer’s reward function, and $\gamma$ is the discount factor.
For the considered anti-jamming MG, we assume that the actions of both the legitimate user and the intelligent jammer are allowed to change among hops. Since the spectrum waterfall defined in Chen and Wen (2016) contains time, frequency and power domain information, it could be used to represent the complex spectrum state and provide enough information for either anti-jamming or jamming decision-making. Therefore, the environment state of the $n$-th hop could be represented as
where ${\bf{s}}_{k}$ denotes the spectrum vector during the $k$-th time slot, $N_{T}$ denotes the length of historical data. The state ${\bf{S}}_{n}\in\mathcal{S}$ is a two-dimensional matrix of size $N_{F}\times N_{T}$, with each row corresponding to the channel power measurements during a given time slot. The thermodynamic chart of the matrix ${\bf{S}}_{n}$ is referred to as the spectrum waterfall Liu et al. (2018), which captures both frequency and temporal information, and all possible observed spectrum waterfalls constitute the set of environmental states. Fig. 4 provides a detailed example of the environmental state in the anti-jamming scenario under consideration. The user takes ${\bf{S}}_{n}$ as the input for anti-jamming decision-making. Similarly, the DRL-based intelligent jammer also takes the observed spectrum waterfall ${\bf{S}}_{n}^{j}$ as the input for its jamming decisions. For convenience, ${\bf{S}}_{n}^{j}={\bf{S}}_{n}$ is used in what follows.
Figure 4: A detailed example of the environmental state.
Let $\pi(\cdot)$ and $\mu(\cdot)$ denote the policies of the legitimate user and intelligent jammer, respectively. During the $n$-th hop, the legitimate user executes an anti-jamming action $a_{n}^{u}\in\mathcal{F}$ based on the policy $\pi({\bf{S}}_{n})$, while the intelligent jammer performs a jamming action $a_{n}^{j}\in\mathcal{F}$ according to the policy $\mu({\bf{S}}_{n})$. Then, the state ${\bf{S}}_{n}$ transits to the next state ${\bf{S}}_{n+1}$ with probability $\mathcal{P}({\bf{S}}_{n+1}|{\bf{S}}_{n},a_{n}^{u},a_{n}^{j})$, and the environment provides immediate rewards $r_{n}^{u}=\mathcal{R}^{u}({\bf{S}}_{n},a_{n}^{u},a_{n}^{j})$ and $r_{n}^{j}=\mathcal{R}^{j}({\bf{S}}_{n},a_{n}^{u},a_{n}^{j})$ to the legitimate user and the intelligent jammer, respectively.
At the end of each hop, the receiver responds to the transmitter with a feedback (i.e., ACK/NACK) through the control link. Specifically, if $\beta(k,f_{k}^{u})\geq\beta_{\text{th}},\ \forall k=(n-1)N_{h}+1,\ldots,nN_{h}-1,nN_{h}$ holds during the $n$-th hop, the receiver transmits an ACK signal to the transmitter. Otherwise, an NACK signal is transmitted. Motivated by Yuan et al. (2024), we assume that the legitimate user’s agent is located at the receiver. The agent could determine which channel to access and transmit the channel decision message $a_{n}^{u}$ to the transmitter at the beginning of the $n$-th hop through the control link. Then it takes the minimum SINR during the $n$-th hop as the immediate reward for action $a_{n}^{u}$, namely,
Whereas the intelligent jammer has a completely contrasting objective. Although the jammer typically lacks knowledge of the received SINR and demodulation threshold of the legitimate user, the ACK/NACK feedback transmitted from the legitimate receiver to the legitimate transmitter could reflect whether the legitimate signal transmitted during a hop has been successfully received Erpek et al. (2019). It is assumed that the worst jammer could accurately assess the effectiveness of jamming attacks, and the reward for the intelligent jammer’s action $a_{n}^{j}$ is determined by the intercepted ACK/NACK signal from the control link during the $n$-th hop, namely,
@@ 2.2 Markov Game Model
In the process of the anti-jamming game, the legitimate user aims to maximize its cumulative reward $R^{u}({\bf{S}}_{n})$, i.e.,
@@
Meanwhile, the intelligent jammer aims to maximize its cumulative reward $R^{j}({\bf{S}}_{n})$, i.e.,
@@
(BR in the anti-jamming MG) Given the policy of the intelligent jammer $\mu$, the policy $\pi^{b}$ of the user is defined as the BR policy if there exists no alternative policy that could yield a higher cumulative reward, formally expressed as
@@
Conversely, when the user employs the policy $\pi$, the BR policy of the intelligent jammer, denoted as $\mu^{b}$, must satisfy
@@
(NE in the anti-jamming MG) The NE is defined as a pair of policies $(\pi^{*},\mu^{*})$, where $\pi^{*}$ and $\mu^{*}$ are mutual best responses. Formally, this could be expressed as
@@
According to Nash (1951), the NE of the proposed MG always exists and is equivalent to the minimax solution of the game, i.e.,
@@
The NE delineates the maximum return that a legitimate user could achieve when facing a formidable opponent. It is particularly meaningful when the jammer is capable of adapting its policy in response to the user’s actions. When the legitimate user adopts the policy $\pi^{*}$, its return is guaranteed to be at least equal to the Nash equilibrium. While, if the legitimate user switches to an alternative anti-jamming policy, its return might fall below $R^{u}({\bf{S}}_{n};\pi^{*},\mu^{*})$.
Minimax Q learning Wang et al. (2011) could be employed to determine the NE for MGs without any prior knowledge of the environment dynamics. Additionally, the legitimate user could achieve a beyond NE performance by employing an opponent modeling based DQN Li et al. (2023c). However, these methods assume that the legitimate user could access the actions selected by the intelligent jammer at the beginning of each hop, which is generally impractical. Meanwhile, these DRL-based methods necessitate extensive training episodes in non-stationary environments, where the strategies of both the legitimate user and the intelligent jammer could continuously evolve over time. To address these issues, we propose a fast adaptive anti-jamming channel access approach to find a policy beyond NE with fewer training episodes in the following sections.